OpenStudy (moongazer):

how do you add minutes to time ? for example: If the current time is 1:30 pm, what is the time after 145 minutes. or If the current time is 1:30 pm, what is the time after 29 hours. and also the subtraction like 90 mins. before etc. Thanks :)

jimthompson5910 (jim_thompson5910):

I would convert 145 minutes to hours, minutes first 145/60 = 2 remainder 25 So 145 minutes = 2 hours, 25 minutes so 1:30 2:25 ----- 3:55 which would mean that it is 3:55 pm

jimthompson5910 (jim_thompson5910):

24 hrs = 1 day 29 hrs = 24+5 hrs 29 hrs = 1 day, 5 hrs so if you add 29 hrs to 1:30, then you're really just adding 5 hrs because you add that full day (which means you go from 1:30 pm on fri to 1:30 pm on sat) then the 5 additional hours